OpenStudy (anonymous):

Evaluate 9x^2^-y-2 for x = –3 and y = 2. A. 324 B.20 1/4 C.9(-6)^0 D.1/144

OpenStudy (pfenn1):

What do you get when you plug in -3 for x and 2 for y in the expression?

OpenStudy (anonymous):

how exactly do you plug it in ?

OpenStudy (pfenn1):

Is this the expression?\[9x^{2^{(-y-2)}}\]

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

without the parenthesis

OpenStudy (anonymous):

answer ?

OpenStudy (pfenn1):

what do you mean without the parenthesis? isn't the 2 raised to the power -y-2?

OpenStudy (anonymous):

oh yeah

OpenStudy (pfenn1):

\[9(-3)^{2^{(-4)}}\]Now we know that \[2^{-4}=\frac{1}{2^4}=\frac{1}{16}\]So now we have \[9(-3)^\frac{1}{16}\] Remember that \[\large x^{\frac ab}=\sqrt[b]{x^a}\]So can you take it from here.

OpenStudy (pfenn1):

@ZachKetchup, you have posted this questions at least 3 or 4 times. Do you not understand how to find the solution? I did almost all the steps for you. Mostly when you are dealing with exponents, there are a bunch of rules that are handy to remember.\[x^0=1\]\[x^{-a}=\frac{1}{a}\]\[x^ax^b=x^{a+b}\]\[(x^a)^b=x^{ab}\]\[\frac{x^a}{x^b}=x^{a-b}\]\[\large x^{\frac ab}=\sqrt[b]{x^a}\]